OpenStudy (anonymous):

How would I calculate something like sin(1) without a calculator?

zepdrix (zepdrix):

You wouldn't. \(\Large\rm \sin(1)\) is as simple as it's going to get. If you had been given something like \(\Large\rm \sin(70^o)\) or \(\Large\rm \sin(15^o)\) then you apply some tricks to relate them to your special angles 30, 45 and such. But I don't think we can do anything with sin(1). Maybe there are some fancy tricks that I'm not aware of though :P

OpenStudy (jdoe0001):

did you mean \(\bf sin^{-1}(1)\quad ?\)

OpenStudy (anonymous):

no, what Im working on is autonomous differential equations and isoclines. So, Im given the differential equation y'=sin(y) and Im supposed to graph a small directional field for it. Now, my book tells me that isoclines for autonomous differential equations are all horizontal. So I know that along each horizontal line, there will be the same slope for each t value.

OpenStudy (anonymous):

The part Im struggling with is figuring out the slope. So, if I understand correctly, I should just take y'=sin(y) , and for each y, say, 2, the slope should be sin(2). But, I have no idea how to figure out what that looks like without a calculator.

OpenStudy (marissalovescats):

My advice is: Memorize the unit circle

zepdrix (zepdrix):

Hmm I'm not sure why you're setting y' and y to be the same values. Your y' ( the slope of the function on the right ) can never be anything except 0, yah? That's what they're telling you in the book. y'=0 everywhere because the slope is horizontal. So I think you want to do something more like this:\[\Large\rm 0=\sin(y)\]And then you would plug 0 in for your `t`'s on the right as well, is that how the process works? I haven't done this in a while. Anyway, you have `t`'s on the right, so you get that the whole region is producing horizontal isoclines ....... or something... is this helpful maybe? blahhhhhhhh

zepdrix (zepdrix):

you have NO* `t`'s on the right* sorry typo

OpenStudy (anonymous):

Hehe, you've given me some things to think about, I appreciate it.